# Improving Assembly Code Performance with Large Language Models via Reinforcement Learning

### v1

@@ C Assembly
We describe our dataset construction approach in Section 3.2. Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation. The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table 1.
@@
For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3. All test cases are withheld from the model. The model is instructed to generate only the optimized x86-64 assembly code. We show the prompt template in Section A.2.
@@
We evaluate each model using both correctness and performance metrics. Compile pass is the percentage of problems for which the generated assembly compiles to binary executable successfully, while test pass is the percentage of problems where the compiled code passes all test cases. For a given problem, any single failed test case is considered a failure for the test pass metric. Both metrics are computed across the entire validation set. For performance, we measure the relative speedup over the gcc -O3 baseline. As defined in Section 3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline. This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain. We report the 25th, 50th (median), and 75th percentiles of speedup to capture distributional behavior, along with the average speedup over the entire evaluation set.
@@
We evaluate 21 state-of-the-art language models spanning a diverse range of architectures. Our benchmark includes frontier proprietary models such as gpt-4o [64], o4-mini, gemini-2.0-flash-001 [65], and claude-3.7-sonnet, as well as open-source families such as Llama [66], DeepSeek [67], and Qwen [25]. In addition, we include models distilled from DeepSeek-R1 [68] based on Qwen and Llama. Finally, we evaluate recent compiler foundation models [39, 40] that are pre-trained on assembly code, building upon Code Llama and designed specifically for compiler-related tasks. All open-source models are instruction-tuned.
@@
To ensure an accurate performance evaluation, we use hyperfine [69], a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions. For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
@@
We implement our customized reinforcement learning reward functions within the verl framework [70], which enables fine-tuning of LLMs using Proximal Policy Optimization (PPO). As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section 3.3. This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
@@
Among all evaluated models (see Table 2), we select Qwen2.5-Coder-7B-Instruct for training due to its strongest correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality. Training is performed on a single node with four A100 GPUs. Full hyperparameter settings are provided in Table A1.
@@ 5.1 Evaluation of Different Models
Table 2: Comparison of LLMs on our assembly optimization benchmark. We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline. All open-source models are instruction-tuned. We evaluate general-purpose foundational models, compiler-specific foundation models, and our PPO-trained model, which improves average speedup from 1.10× to 1.47×.
Table 2 presents results across evaluated models. Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates. Among baseline models, claude-3.7-sonnet and DeepSeek-V3 perform best, achieving test pass rates above 40% and average speedups of 1.22× and 1.21×, respectively. Notably, some models such as DeepSeek-R1 fail to generate any valid assembly, and o4-mini achieves only 4.5% test pass. These results underscore the difficulty of the task and motivate the need for a task-specific approach.
Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler intermediate representations. Among them, llm-compiler-13b demonstrates strong performance in both correctness and speedup. In contrast, the fine-tuned variants (-ftd) perform poorly, likely because they are adapted for tasks such as disassembling x86-64 and ARM assembly into LLVM-IR, rather than optimizing assembly code for execution performance.
We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong compile pass rate (79.0%) and highest test pass rate (61.0%) among models. After PPO fine-tuning, it achieves 96.0% on both metrics and increases average speedup from 1.10× to 1.47×. Notably, it is the only model to exhibit meaningful speedup even at the 25th percentile, and it outperforms all other models across all evaluation metrics, including correctness, average speedup, and speedup percentiles.
Method Compile Pass Test Pass Avg. Speedup
Base Model 79.0% 61.0% 1.10x
RL w/ CGS 95.5% 94.5% 1.38×
RL w/ SO 96.0% 96.0% 1.47×
Table 3: Ablation study comparing reward function variants. CGS provides intermediate reward shaping, while SO uses a sparse and terminal signal.
5.2 Ablation Study of Reward Function Design
We evaluate two reward designs for RL: Correctness-Guided Speedup (CGS) and Speedup-Only (SO). CGS penalizes compilation failures, rewards partial correctness, and scales final reward by speedup once all tests pass. SO uses speedup as the sole reward, but only when all tests pass.
As shown in Table 3, both variants achieve high compile pass rates and test pass rates, but SO yields better performance. Removing intermediate shaping appears to help the model focus on terminal objectives. We also tried varying the CGS scaling factor $\alpha$ (5 or 10) and found that it has a negligible effect.
These results indicate that sparse, terminal rewards (SO) are more effective in this setting. Since the base model already reaches 61.0% test pass, correctness is not the bottleneck; optimizing directly for speedup offers a stronger training signal.
Method Compile Pass Test Pass Avg. Speedup
w/ O3 w/o O3 w/ O3 w/o O3 w/ O3 w/o O3
DeepSeek-V3 94.0% 25.5% 43.0% 4.5% 1.21× 1.02×
claude-3.7-sonnet 94.5% 53.0% 58.5% 16.0% 1.22× 1.07×
llm-compiler-7b 55.0% 12.0% 54.0% 0.0% 1.09× 1.00×
llm-comiler-13b 60.5% 2.0% 59.5% 0.5% 1.34× 1.00×
Qwen2.5-Coder-7B 79.0% 0.0% 61.0% 0.0% 1.10× 1.00×
Qwen2.5-Coder-7B-PPO 96.0% 0.0% 96.0% 0.0% 1.47× 1.00×
Table 4: Ablation study on the impact of including gcc -O3 baseline assembly in the prompt. Each metric is reported with and without access to the baseline assembly generated by the compiler.
5.3 Can LLMs Directly Compile Programs without Baseline Assembly?
In our main evaluation, we always include the baseline assembly generated by gcc -O3 in the prompt. While this baseline offers a strong starting point for further optimization, it may also bias the model toward replicating patterns from the compiler’s output. In this subsection, we investigate a more challenging setup: Can large language models directly compile C code into performant assembly without relying on the compiler-generated baseline?
We compare two settings in evaluation: (1) providing both the C source and the gcc -O3 assembly (default, “w/ O3” in Table 4), and (2) providing only the C source (“w/o O3”). For each model, we report compile success rate, test pass rate, and average speedup.
Table 4 shows that removing the baseline assembly leads to severe degradation. For instance, Qwen2.5-Coder-7B-PPO drops from 96.0% correctness and 1.47× speedup to 0.0% and 1.00×, respectively. Even strong models like Claude-3.7-sonnet suffer substantial declines.
These results suggest that direct compilation from C to optimized assembly remains challenging for current LLMs. The compiler output provides a reliable reference for LLMs. This supports our framework design: using gcc -O3 as an effective starting point for reinforcement learning. While future work may explore direct generation from C, we expect it to be substantially more challenging due to the absence of compiler guidance.
5.4 Case Study
@@
Figure 2 presents a representative example where large language models (LLMs), including gpt-4o and claude-3.7-sonnet, discover an optimization that outperforms a state-of-the-art compiler. The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit. The assembly code produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches to compute the result.
In contrast, the LLM generates a significantly more concise and efficient implementation that replaces the entire loop with a single popcnt instruction. This instruction, supported by modern x86-64 architectures, performs the same computation in one operation, thereby reducing both instruction count and runtime overhead.
Such a transformation is beyond the reach of gcc -O3, which applies a predetermined sequence of rule-based optimization passes and does not conduct semantic-level rewrites of this kind. In this case, the language model is able to synthesize functionally equivalent code that exploits hardware-level instructions not utilized by the compiler. This demonstrates the potential of language models to optimize assembly by exploring a broader space of semantics-preserving program transformations.
@@ C Code
Figure 2: Case study comparing the C code, baseline assembly produced by gcc -O3, and optimized assembly generated by the LLM. The model successfully replaces the loop with the specialized hardware instruction popcnt, resulting in a significantly more concise implementation.
6 Discussion
Limitations.
A key limitation of our approach is the absence of formal correctness guarantees. Although we validate generated programs using input-output test cases, such testing is inherently incomplete and may overlook edge cases. Consequently, unlike fully verified systems such as Stoke [11], LLM-generated assembly may produce wrong results or exhibit undefined behavior. This limitation reflects a broader challenge in programming languages: verifying the semantic equivalence between two arbitrary programs is undecidable in the general case.
Another limitation lies in the inherent randomness of performance measurement on real hardware. Although we mitigate noise through repeated measurements, low-level hardware fluctuations can still introduce variability into speedup estimates. Such nondeterminism is difficult to eliminate entirely in real-world settings. While prior work has adopted simulator-based evaluation [6], simulators may fail to faithfully capture the actual hardware performance.
Finally, the observed performance gains may not generalize across machines. The model may implicitly learn and exploit hardware-specific characteristics like cache size. As a result, a model trained on one machine may not retain its effectiveness when deployed on a different machine.
Future Work.
While we use Proximal Policy Optimization (PPO), future work may explore alternative reinforcement learning algorithms such as GRPO [71]. Expanding to larger and more diverse datasets, particularly those involving performance-critical code beyond competitive programming, would make the setting more realistic and applicable. Combining reinforcement learning with supervised fine-tuning may also be beneficial, although it remains unclear whether training on gcc -O3 outputs would provide additional gains. Another direction is to incorporate an interactive refinement loop, where the model iteratively updates its output using feedback from errors or performance measurements. Finally, extending our approach from x86-64 to other architectures such as MIPS, ARM, or GPU programming could broaden its applicability and impact.
7 Conclusion
We explore the use of large language models (LLMs) for optimizing assembly code, a setting where fine-grained control over execution enables performance improvements that are difficult to express in high-level languages. While traditional compilers rely on fixed rule-based transformations, they face fundamental limitations due to the complexity of the optimization space. To address this, we apply reinforcement learning to fine-tune LLMs with Proximal Policy Optimization (PPO), using a reward function based on correctness and speedup over the gcc -O3 baseline. To support this effort, we introduce a benchmark of 8,072 real-world C programs with compiler-generated baseline assembly and test cases. Our resulting model, Qwen2.5-Coder-7B-PPO, achieves the highest compile and test pass rates (96.0%) and the best average speedup (1.47×), outperforming all 20 other models evaluated across all metrics. These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
Appendix A Appendix
A.1 Training Configurations
@@
Table A1: Key training configurations for VERL PPO fine-tuning.
A.2 Prompt Template

### v2

@@ 4 Experimental Setup
We describe our dataset construction approach in Section˜3.2. Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation. The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table˜1, and additional analysis below.
Test Coverage.
Our dataset includes test cases for every program. Rather than relying directly on the original submissions, we re-run each program on its inputs to generate correct outputs, thereby fixing errors in prior datasets. The resulting test suites of our evaluation dataset achieve an average of 96.2% line coverage and 87.3% branch coverage, demonstrating high test quality.
@@ C Assembly
Speedup by Compilers.
We quantify compiler optimizations by comparing gcc -O0 with gcc -O3 on the evaluation dataset and observe a mean speedup of 2.65×. This demonstrates the substantial effect of compiler optimizations and confirms that performance improvements in our dataset are measurable. Building on this baseline, we investigate whether LLMs can further enhance performance beyond the 2.65× speedup provided by the compiler.
@@
For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3. Test cases are withheld from the model. The model is instructed to generate the optimized assembly code. We show the prompt template in Appendix A.3.
@@
We evaluate each model using both correctness and performance metrics. Compile pass is the percentage of problems for which the generated assembly compiles to binary executable successfully, while test pass is the percentage of problems where the compiled code passes all test cases. For a given problem, any single failed test case is considered a failure for the test pass metric. Both metrics are computed across the entire validation set. For performance, we measure the relative speedup over the gcc -O3 baseline. As defined in Section˜3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline. This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain. We report the 25th, 50th (median), and 75th percentiles of speedup to capture distributional behavior, along with the average speedup over the entire evaluation set.
@@
We evaluate 23 state-of-the-art language models spanning a diverse range of architectures. Our benchmark includes frontier proprietary models such as gpt-4o (Achiam et al., 2023), o4-mini, gemini-2.0-flash-001 (Team et al., 2023), and claude-3.7-sonnet, as well as open-source families such as Llama (Touvron et al., 2023), DeepSeek (Liu et al., 2024a), and Qwen (Hui et al., 2024). In addition, we include models distilled from DeepSeek-R1 (Guo et al., 2025) based on Qwen and Llama. Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024; 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table˜2). All open-source models are instruction-tuned.
@@
To ensure an accurate performance evaluation, we use hyperfine (hyp, 2025), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions. For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
@@
We implement our customized reinforcement learning reward functions within the verl framework (Sheng et al., 2024), which enables fine-tuning of LLMs using PPO and GRPO. As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section˜3.3. This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
@@
Among all evaluated models (see Table˜2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality. Training is performed on a single node with four A100 GPUs. Full hyperparameter settings are provided in Appendix A1.
@@ 5.1 Evaluation of Different Models
Table 2: Comparison of LLMs on our assembly optimization benchmark. We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline. All open-source models are instruction-tuned.
Main Results.
Table˜2 reports results across evaluated models. Most perform poorly on this task, with only a few demonstrating effectiveness as superoptimizers. Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates. Among all models, claude-opus-4 and claude-sonnet-4 perform best, achieving average speedups of 1.43× and 1.30×, respectively. Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler IRs. Among them, llm-compiler-13b achieves a notable 1.34× speedup, whereas the fine-tuned variants (-ftd) perform poorly, likely because they were adapted for different tasks such as disassembling x86-64 into LLVM IR. These results suggest that while superoptimization is inherently difficult, some LLMs can be effective superoptimizers.
Failure Modes.
Interestingly, models that are expected to achieve strong results perform (e.g., DeepSeek-R1, GPT-4o) perform poorly on the task of superoptimization, motivating our analysis of their failure modes. We find that DeepSeek-R1 consistently fails to generate valid assembly code, resulting in a 0% compilation rate. DeepSeek-R1 often produces verbose analysis instead of executable code, spending its entire output length on reasoning about instruction semantics and potential optimizations without actually implementing them.
We further analyze the failure modes of GPT-4o, which achieves a high compilation rate (81.0%) but exhibits poor correctness (only 5.0% test pass rate). The primary correctness issues are as follows: (1) missing critical directives and safety setup, such as stack canary initialization and .cfi_* metadata, which often lead to runtime crashes; (2) incorrect function call conventions, where repeated system calls like scanf are made without proper argument setup, causing undefined behavior; (3) semantic errors in core computations, including incorrect pointer usage or altered algorithm logic, which produce wrong outputs even when the code runs; and (4) over-simplified stack or register management, resulting in memory errors or invalid control flow. In summary, GPT-4o tends to sacrifice correctness in pursuit of optimization: it generates syntactically valid assembly but frequently violates low-level conventions necessary for correct and reliable execution.
5.2 Effectiveness of RL Training
Improvement. Table˜3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements. We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong test pass rate (61.4%) among models. After RL training with PPO, the fine-tuned model SuperCoder attains 95.0% correctness and improves average speedup from 1.10× to 1.46×. Its speedup percentiles are 1.17 ± 0.03× (25th), 1.35 ± 0.04× (50th), and 1.64 ± 0.08× (75th) respectively, outperforming the majority of evaluated models.
PPO versus GRPO. We evaluate both PPO-trained and GRPO-trained models and find their performance to be nearly identical. SuperCoder trained with GRPO attains 94.7 ± 0.6% correctness and 1.44 ± 0.07× average speedup, which is comparable to SuperCoder trained with PPO (95.0 ± 0.0% correctness and 1.46 ± 0.12× average speedup).
Model Compile Pass Test Pass Average Speedup
[…]
Table 3: Performance of the base model and the models fine-tuned with RL. Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
5.3 Analysis of Learned Program Transformations
To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by our PPO-trained model SuperCoder. We categorize the learned transformations into four types: (1) code layout and instruction scheduling, which account for 98.5% of the changes and include basic block placement, instruction alignment, and reordering to improve cache behavior and hide latencies; (2) register allocation, representing 16% of the changes, where the model selects alternative registers; (3) control flow optimization (1.5%); and (4) instruction selection (1.0%). Since a single problem’s transformation may belong to multiple categories, the percentages do not sum to 100%. Unfortunately, there is currently no automated tool that can further reliably explain performance speedups at the assembly level. Automated analysis and explanation of performance differences in assembly code remains a challenging and unsolved problem.
The observed gains and our analysis suggest that LLMs can produce reasonable code transformations that meaningfully improve assembly performance. These results indicate that LLMs can uncover nuanced optimization opportunities beyond the reach of traditional compiler heuristics, suggesting that compilers themselves still have room for improvement.
5.4 Comparison with Simple Random Perturbations
To assess whether the improvements achieved by LLMs stem from systematic strategies learned during training rather than simple random variation in assembly programs, we compare against a baseline that applies simple random perturbations to assembly code. While such perturbations do not necessarily improve performance, running them multiple times and selecting the fastest version allows us to compute a speedup. These perturbations preserve correctness but can still influence performance, and we report speedup using the minimum observed runtime across 10 runs.
To randomly perturb the assembly, we compile the binaries with options that randomize both the stack and program load addresses. This alters code layout in memory, which can affect performance but not correctness. Speedup is measured as the ratio of the fastest run to the original runtime over 10 randomized runs. Repeating the experiment five times yields a speedup of 1.19 ± 0.01× (95% CI), significantly lower than 1.46 ± 0.12× achieved by SuperCoder. This indicates that SuperCoder systematically exploits performance opportunities rather than relying on simple random perturbations, particularly since it produces each result in a single shot without retries or iterative refinement.
6 Discussion
Alternative Reward Function Design.
Besides the reward function presented in Section˜3.3, we also evaluate an alternative design. The original design assigns zero reward whenever any test fails. In contrast, the alternative assigns (i) a reward of 1-1 if the program fails to compile, (ii) a partial reward equal to the fraction of passed tests if only some tests succeed, and (iii) a reward of $1+\alpha\cdot\text{speedup}$ once all tests pass. Training the base model with this design yields an average speedup of 1.38×. Varying the scaling factor (5 or 10) has a negligible effect, and the result remains worse than the 1.46× achieved by SuperCoder with the original reward. This suggests that directly optimizing for the terminal speedup reward is more effective, since the alternative design allows models to avoid penalties by merely compiling or even gain positive reward from partially passing tests, which does not translate into actual speedup.
Direct Compilation from C.
We also examine a more challenging setting where gcc -O0 assembly is not provided. Instead, LLMs receive only the C source code and are asked to generate assembly directly. This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it. Similar degradation occurs for other models such as llm-compiler-13b and Claude models. These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
Case Study.
@@ 6 Discussion
We illustrate in Appendix A.1 a representative example where an LLM discovers an optimization beyond the reach of a state-of-the-art compiler. The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit. The assembly produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches. In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation and reducing both instruction count and runtime overhead.
Limitation and Potential Directions.
A limitation of our approach is the lack of formal correctness guarantees. Although we validate generated programs with input–output test cases, testing is inherently incomplete and may miss corner cases even with high coverage. However, to our knowledge, no existing verifier can generally handle our assembly programs with loops. We hope this work motivates future research in the formal methods communities toward scalable verification techniques. The dataset we introduce could serve as a meaningful benchmark for such research.
Compared with traditional compiler optimization, reinforcement learning introduces substantial computational overhead during training. This cost, however, is incurred only once: once trained, the model can generate optimized assembly efficiently with low latency.
One promising direction for future work is to incorporate an interactive refinement loop, where the model iteratively improves its output based on feedback from errors or performance measurements. In addition, most superoptimization research has centered on x86-64, and our work follows this focus. Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability and impact. Such extensions would mainly require adapting the data collection and testing infrastructure, while keeping the core methodology unchanged.
7 Conclusion
We investigated whether LLMs can act as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. To support this study, we introduced the first large-scale benchmark of 8,072 assembly programs. Evaluating 23 models revealed the difficulty of the task, with most failing to achieve meaningful gains. By fine-tuning with reinforcement learning, our model SuperCoder improved both correctness and performance, reaching 95.0% test pass rate and an average speedup of 1.46× over gcc -O3. These results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly code, opening new opportunities for performance optimization beyond compiler heuristics.
LLM Usage
Large Language Models (LLMs) are the primary subject of study in this paper. In addition, we used LLMs as a general-purpose writing assistant to polish the presentation and improve readability.
Appendix A Appendix
A.1 Case Study
@@ C Code
Figure A1: Case study comparing the C code, baseline assembly produced by gcc -O3, and optimized assembly generated by Claude-Opus-4. The model successfully replaces the loop with the specialized hardware instruction popcnt, resulting in a significantly more concise implementation.
A.2 Training Configurations
@@
Table A1: Key training configurations for PPO fine-tuning.
A.3 Prompt Template